OpenStudy (anonymous):

Find a polar equation for the curve represented by the given Cartesian equation. x + y = 2

OpenStudy (perl):

you can replace x with r cost theta, and y with r sin theta

OpenStudy (anonymous):

so rcost + rsint = 2 But I'm confused on what to do from then

OpenStudy (perl):

factor out r

OpenStudy (perl):

r = 2 / ( cos t + sin t )

OpenStudy (anonymous):

Then r^2(cost+sint)^2 = 4 r^2 (1 + 2sintcost) = 4 ?

OpenStudy (perl):

that wont simplify it

OpenStudy (perl):

once you have a function r = f(theta), you are done

OpenStudy (anonymous):

ohhh I see.

OpenStudy (anonymous):

Thank u much!

OpenStudy (perl):

the other direction is harder, converting polar equation to cartesian

OpenStudy (perl):

sure :)

OpenStudy (perl):

here is a fun problem find the polar equation for the cartesian equation y = x

OpenStudy (anonymous):

y - x = 0 y^2 + x^2 -2xy = 0 1 -2(rcost * rsint) = 0 -2(rcost * rsint) = -1 2(rcost * rsint) = 1 (rcost * rsint) = 1/2 r(cost*sint) / 1/2 r = 1/2sintcost r = 1/sin2t Is that right?

OpenStudy (perl):

x^2 + y^2 does not equal to 1

OpenStudy (anonymous):

oh it's equal to r^2 right. let me try again

OpenStudy (perl):

wait

OpenStudy (perl):

why not just substitute y = x , right away

OpenStudy (anonymous):

So x^2 + x^2 -2x^2 = 0 wouldn't that just give 0 = 0?

OpenStudy (perl):

y = x r sin theta = r cos theta

OpenStudy (anonymous):

So sin theta = cos theta sin t / cos t = 1 tan t = 1

OpenStudy (perl):

looking better

OpenStudy (anonymous):

And that makes sense because there should be no r in the function, since it's just in one direction, which is when tangent is 1 and that's at y =x.

OpenStudy (perl):

right, you can keep solving theta = arctan (1) = pi/4

OpenStudy (perl):

so you have polar coordinates : ( r , pi/4) r can vary

OpenStudy (anonymous):

I see. You made this very clear to me. Thanks again :)

OpenStudy (perl):

:)